OpenStudy (anonymous):

Marc is outlining a square play area around a swing set in their backyard. He lays 60 feet of wooden planks around the entire outline. To the nearest tenth, what is the diagonal distance from one corner of the play area to the opposite corner? A. 21.2 ft B. 30.0 ft C. 42.4 ft D. 84.9 ft

OpenStudy (anonymous):

The answers A because you know that the outline is a square and your given a total of 60 feet of wooden planks. So each side of the play area is 15 ft each and you have to solve for the diagonal distance. Therefore, using pythagorean theorem: a^2 + b^2 = c^2 15^2 + 15^2 = c^2 c = 21.2

OpenStudy (anonymous):

thank you